OpenStudy (mathmath333):

At a game of billiards, A can give B 15 points in 60 and A can give C to 20 points in 60. How many points can B give C in a game of 90?

OpenStudy (nurali):

"A can give 15 points to B in 60". In a game of 60, A starts with 0 points where B starts with 15 points "A can give 20 points to C in 60". In a game of 60, A starts with 0 points where C starts with 20 points ie, in a game of 60-15-45, B can give C 20-15=5 points ie, in a game of 90, B can give C 5×2=10 points

OpenStudy (mathmath333):

i never played pool , i m having difficulty to understand this.

OpenStudy (rational):

A/B = 60/15 A/C = 60/20 B/C = ?

OpenStudy (rational):

forget about billards the question is about ratios

OpenStudy (mathmath333):

yes but i m bit curious here, i believe in trying to visualize the imagination..

OpenStudy (mathmath333):

this question is from time , speed and distance and , how is it related to that

OpenStudy (ikram002p):

So we have to interrupt 15 in 20 as a game of 90 ?

OpenStudy (mathmath333):

i will go and play billiards tommorrow , to understand better :D